OpenStudy (anonymous):

find the volume of the solid using shell method, region of y=sqrt2, x=y^2 , x=0. about the x-axis

OpenStudy (anonymous):

Ok which function is on top?

OpenStudy (anonymous):

y=sqrt2, bounded on left by y -axis and below by x=y^2

OpenStudy (anonymous):

just confirming it is y=\[y=\sqrt(2)\] x= x^2 right functions?

OpenStudy (anonymous):

-- correction x=y^2

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Ok so x=y^2 is a side way parabola facing in the positive x direction(right) y=sqrt(2) is just a horizonal striaght line We need to find intersection of of lines

OpenStudy (anonymous):

y=sqrt2 right?

OpenStudy (anonymous):

The intersection point will have y=sqrt(2)

OpenStudy (anonymous):

when x = 2

OpenStudy (anonymous):

what is x intersection is 2

OpenStudy (anonymous):

so (2,sqrt(2)

OpenStudy (anonymous):

good so far?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

So the region we are rotating has these points(0,0)-- origin (0,sqrt(2)top left corner, (2,sqrt(2) intersection

OpenStudy (anonymous):

any questions?

OpenStudy (anonymous):

so far good, i have a drawing set up

OpenStudy (anonymous):

So we are going to use shell method which is just add up bunch off cylinder

OpenStudy (anonymous):

Volume of cylinder is pi*radius^2*h

OpenStudy (anonymous):

got ya

OpenStudy (anonymous):

So if you look your drawing(may be draw horizonal rectangle) you will see that the y value of function represent radius

OpenStudy (anonymous):

and the function x=y^2 represent h

OpenStudy (anonymous):

yea from the integral of 0 to sqrt2 right?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

\[2\pi\int\limits(y*y^2 dy(\]

OpenStudy (anonymous):

you got the limit of the integration right

OpenStudy (anonymous):

so just integrate that? with the height as y^2?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

oooh thank you, i kept thinking of subtracting sqrt2 from the height wasn't really paying attention that this was respect to y. Thank you for all your help:)